# Adapted optimal transport between Gaussian processes in discrete time

### v1

@@ 7. Conclusion
In this paper we derived an expression for the adapted Wasserstein distance between Gaussian distributions and characterized the optimal couplings. Apart from a deeper study of the geometry of adapted Wasserstein distance, we may consider the adapted Wasserstein barycenter and entropic regularization of (2.6) with Gaussian marginals. Motivated by the Wasserstein setting [1, 17], it seems natural to conjecture that the solutions to both problems are Gaussian. It is also of interest to use the explicit formula of $d_{\mathrm{ABW}}$ to study, in the spirit of [16], distributionally robust optimization problems under Gaussian dynamics. Finally, our results may be useful for studying the adapted Wasserstein distance between continuous time Gaussian processes which are not necessarily semimartingales.

### v4

@@ 7. Conclusion
In this paper we derived explicitly the adapted Wasserstein distance between non-degenerate Gaussian distributions and characterized the optimal couplings. We also studied some properties of the resulting adapted Bures-Wasserstein distance between covariance matrices. Our results, as well as the extensions in [1], suggest many directions for future research. Apart from a deeper study of the adapted Bures-Wasserstein distance (including the degenerate case), we may consider the adapted Wasserstein barycenter introduced in [2]. Motivated by the Wasserstein setting [3], it seems natural to conjecture that when the marginals are Gaussian, there exists an adapted Wasserstein barycenter which is Gaussian. It is also of interest to use the explicit formula of $d_{\mathrm{ABW}}$ to study, in the spirit of [19], distributionally robust optimization problems, as well as filtering, projection and stochastic control under adapted Wasserstein loss. Finally, our techniques may be useful for studying the adapted Wasserstein distance between continuous-time Gaussian processes which are not necessarily semimartingales.